### v3

# Quantum Thermodynamic Integrability for Canonical and non-Canonical Statistics

Abstract
We extend the Carathéodory principle of the Second Law to quantum thermodynamics with energy levels depending on macroscopic variables, such as volume and magnetic field. This extension introduces the concept of Quantum Thermodynamic Integrability (QTI), offering an alternative foundation for statistical mechanics. QTI is characterized by the path-independence of work and heat within the thermodynamic manifold, which is locally described by energy levels and specific thermodynamic parameters. Within this framework, temperature naturally emerges as an integrating factor, allowing for the derivation of both canonical and non-canonical states from the Entropy Integrable Equations (EIE) based on QTI. Notably, non-canonical states, which become particularly significant outside the thermodynamic limit, reveal the existence of informational correlations in finite-size thermodynamic systems.
Introduction- Statistical mechanics provides the microscopic foundation of thermodynamics, bridging the gap between the macroscopic properties of systems and their underlying microscopic behaviors (Schrödinger, 1952; Hatsopoulos and Keenan, 1965; Gallavotti, 1999; Landau and Lifshitz, 2013). It aims to elucidate thermodynamic phenomena through the statistical behavior of a large number of particles. Despite its profound significance, the foundational principles of statistical mechanics remain subjects of ongoing debate (Khinchin, 1949; Jaynes, 1957a, b; Goldstein et al., 2006; Dong et al., 2007; Landau and Lifshitz, 2013; Gogolin and Eisert, 2016). Key issues include the use of ensembles versus time averages (Khinchin, 1949; Landau and Lifshitz, 2013), the equal a priori probability assumption (Khinchin, 1949; Landau and Lifshitz, 2013), the maximum entropy principle (Jaynes, 1957a, b), and canonical typicality based on quantum entanglement (Goldstein et al., 2006; Popescu et al., 2006; Dong et al., 2007; Gogolin and Eisert, 2016). However, statistical mechanics relies on these foundational assumptions to construct statistical distributions, which lack direct experimental validation. This raises a fundamental inverse question: can the statistical distribution of systems be derived directly from the fundamental principles of thermodynamics, which have been solidly supported by experimental evidence since the era of Watt’s steam engine?
The fundamental principles of thermodynamics were even thought to emerge from the concept of thermodynamic integrability, as originally proposed by Carathéodory (Carathéodory, 1909; Born, 1921; LANDSBERG, 1964; Pogliani and Berberan-Santos, 2000; Ma et al., 2021). Unlike the conventional formulation of the Second Law of Thermodynamics, which relies on the cycle operation of heat engines, Carathéodory’s axiomatization is grounded in classical thermodynamic integrability. This approach is characterized by the path-independence of work, heat and entropy in thermodynamic spaces defined by macroscopic parameters such as volume and magnetic field. In the realm of quantum thermodynamics, internal energy, work, and heat are all described in terms of discrete energy levels and their distributions (Schrödinger, 1952; Quan et al., 2005, 2007). To extend Carathéodory’s approach to quantum thermodynamics, we introduce additional thermodynamic quantities that complement the mechanical variables controlling the energy levels.
In this letter, we align with the concept of quantum thermodynamic integrability (QTI) as an extension of the Carathéodory principle. As a core result, the Entropy Integrable Equations (EIE) derived from QTI provide a comprehensive description of various distributions in statistical mechanics, encompassing both canonical and non-canonical states. At equilibrium with detailed balance, the ratio of the probabilities of two energy levels is independent of any other levels, and then the canonical states with Boltzmann distribution is uniquely determined by QTI. We also show that, as more general solutions of the EIE, non-canonical states arise in finite systems outside the thermodynamic limit, which are of particular interest. These non-canonical states imply the presence of informational correlations and often represent statistical distributions characterized by a superposition of different temperatures. A notable example of a non-canonical state is Hawking radiation emitted from small black holes, which is highly relevant to the information loss paradox (Hawking, 1975, 1976; Parikh and Wilczek, 2000; Zhang et al., 2009; Cai et al., 2016; Dong et al., 2014).
Quantum Thermodynamic Integrability - Consider an $N$-level quantum system with a Hamiltonian $H(\bm{\lambda})=K+V(\bm{\lambda})$, which depends on the macroscopic mechanical variables $\bm{\lambda}=(\lambda_{1},\lambda_{2},\cdots,\lambda_{M})\in\mathbb{R}^{n}$, such as the volume and magnetic field. Here $K$ is the kinetic energy, and $V(\bm{\lambda})$ denotes the potential. The eigenstates $|n(\bm{\lambda})\rangle$ and energy levels $E_{n}=E_{n}(\bm{\lambda})$ are controlled by $\bm{\lambda}$ (sometimes subject to boundary conditions). These energy levels form the energy submanifold (ESM): $\{\bm{E}=(E_{1},E_{2},\cdots,E_{N})\}$. For example, the energy levels of a free particle trapped in a one dimensional box are determined by the width of the box $L$ by $E_{n}=\pi^{2}\hbar^{2}n^{2}/2mL^{2}$.
In the eigenstates $|n(\bm{\lambda})\rangle$ defined by $H(\bm{\lambda})|n(\bm{\lambda})\rangle=E_{n}(\bm{\lambda})|n(\bm{\lambda})\rangle$, the probability distribution $P_{n}=P_{n}(\bm{E})$ forms a vector $\bm{P}\equiv(P_{1},P_{2},\cdots,P_{N})$ that depends on $\bm{E}$. When the potential well is adjusted slowly by $\bm{\lambda}$, the system adiabatically stays in the state $|n(\bm{\lambda})\rangle$, and its work changes as
$\delta W_{n}=-\langle n|\partial V/\partial\bm{\lambda}|n\rangle\cdot\text{d}\bm{\lambda}=-(\partial E_{n}/\partial\bm{\lambda})\cdot\mathrm{d}\bm{\lambda},$ (1)
according to the Feynman-Hellmann theorem (Feynman, 1939). Therefore, the change in the average work reads $\delta W=\bm{X}\cdot\mathrm{d}\bm{\lambda}$ (like $\delta W=P\mathrm{d}V$ for ideal gases), where $\bm{X}\equiv\sum_{n}P_{n}\partial E_{n}/\partial\bm{\lambda}$ is the generalized force corresponding to $\bm{\lambda}$. In a general situation, the variation of work is rewritten as $\delta W=\mathrm{tr}(\rho\mathrm{d}H)$ in terms of the density matrix $\rho$. Thus, a change in the internal energy $U=\mathrm{tr}(\rho H)$ can be separated into the work $\delta W$ and the heat $\delta Q=\mathrm{tr}(H\mathrm{d}\rho)$, accordingly.
To generalize the Carathéodory principle to the quantum realm, we assume that the density matrix for the equilibrium state not only depends on $\bm{\lambda}$, but also on some thermodynamic variables $\xi$ to be considered, namely, $\rho=\rho(\bm{\lambda},\xi)$. In the thermodynamic manifold (TDM): $\{(\bm{E}(\bm{\lambda}),\xi)\}$ parameterized by $\{(\bm{\lambda},\xi)\}$ (see FIG. 1), the heat reads as
$\delta Q=\mathrm{tr}(H\partial_{\mu}\rho)\mathrm{d}\lambda_{\mu}+\mathrm{tr}(H\partial_{\xi}\rho)\mathrm{d}\xi.$
Here, $\partial_{\mu}\equiv\partial/\partial\mu$, $\partial_{\xi}\equiv\partial/\partial\xi$, and the repeated index in a term implies a summation. Obviously, the Carathéodory principle can work well in the parameterized TDM : $\{(\bm{\lambda},\xi)\}$, establishing the so-called quantum thermodynamic integrability (QTI).
Figure 1: Extending from energy submanifold (ESM) to the thermodynamics manifold (TDM) shown with two levels system. (a) In the energy submanifold, the vanishing work integration over a close path implies the quantum thermodynamic integrability (QTI). (b) In the thermodynamics manifold where the ESM is its submanifold the vanishing entropy integration over a close path implies the QTI.
Entropy integrable equations for statistical distribution - The QTI requires that there exist an integrating factor $\beta=\beta(\bm{\lambda},\xi)$ in TDM, such that the thermodynamic entropy change $\delta S\equiv\beta\delta Q$ is a total differential, guaranteed with $\mathrm{d}(\delta S)=0$. This gives an equation in two-form describing the integrability of entropy. We take a special solution $\beta=\xi$, which is the simplest integrating factor. Here $\beta=1/k_{B}T$ can be regarded as the inverse temperature. Consequently, the requirement of QTI reads as
$\mathrm{tr}\left(\partial_{\mu}H\partial_{\nu}\rho-\partial_{\nu}H\partial_{\mu}\rho\right)$ $=0,$ (2)
$\mathrm{tr}\left(H\partial_{\mu}\rho-\beta\partial_{\mu}H\partial_{\beta}\rho\right)$ $=0,$ (3)
where $\partial_{\beta}\equiv\partial/\partial\beta$. It is noteworthy that the QTI for composed system guarantees the universality of the integrating factor $\beta$. For example, a system is separated into two portions $1$ and $2$. Denote the integrating factor of the two portions as $\beta_{1}$ and $\beta_{2}$. Consequently, the thermodynamic entropy for subsystem $i=1,2$ reads $\mathrm{d}S_{i}=\beta_{i}\delta Q_{i}$. The additivity of heat $\delta Q=\delta Q_{1}+\delta Q_{2}$ guarantees the additivity of entropy $\mathrm{d}S=\mathrm{d}S_{1}+\mathrm{d}S_{2}$, when $\beta_{1}=\beta_{2}=\beta$, namely, the integrating factor $\beta$ is uniform for all systems in equilibrium.
Typically, with the diagonal representation of density matrix $\rho=\sum_{n}P_{n}|n\rangle\langle n|$, the probability distribution $P_{n}=P_{n}(\bm{E}(\bm{\lambda}),\beta)$ relies on $\bm{\lambda}$ through $\bm{E}(\bm{\lambda})$. Thus, Eqs. (2) and (3) are reduced into
$\sum_{m<n}\left(\partial_{m}P_{n}-\partial_{n}P_{m}\right)\{E_{m},E_{n}\}_{\mu\nu}$ $=0,$ (4)
$\left(\partial_{\mu}\bm{E}\right)\cdot\left[(\bm{\bm{E}\cdot\nabla})-\beta\partial_{\beta}\right]\bm{P}$ $=0,$ (5)
where $\partial_{n}\equiv\partial/\partial E_{n}$, $\{A,B\}_{\mu\nu}\equiv(\partial_{\mu}A\partial_{\nu}B-\partial_{\nu}A\partial_{\mu}B)$ and $\bm{\nabla}\equiv\left(\partial_{1},\partial_{2},\cdots,\partial_{N}\right)$.
It is observed that the QTI is sufficiently guaranteed by the following equations:
$\partial_{m}P_{n}-\partial_{n}P_{m}$ $=0,$ (6)
$\left[(\bm{\bm{E}\cdot\nabla})-\beta\partial_{\beta}\right]\bm{P}$ $=0.$ (7)
Eqs. (6) and (7), referred to as the Entropy Integrable Equations (EIE), represent our core results. As sketched with two energy levels in FIG. 1, the two EIEs represent the QTI within the ESM (FIG. 1(a)) and TDM (FIG. 1(b)) , respectively.
The universal solution $\bm{P}=\bm{\nabla}F(\beta,\bm{E})$ to Eq. (6) can be expressed in terms of a generating function $F$. The normalization of the probability distribution requires $e\cdot\bm{\nabla}F=1$, for $\bm{e}=(1,1,\cdots,1)$. The function $F(\beta,\bm{E})$ can be understood as the generalized free energy, and Eq. (7) is reduced to
$\bm{E}\cdot\bm{\nabla}F=\partial_{\beta}(\beta F)=F+\beta\partial_{\beta}F.$ (8)
For a detailed derivation of the EIE, please see Sec. I of the Supplementary Materials (SM_, ). Here, the normalization condition guarantees that the distribution is independent of the reference point of energy. In fact, for a uniform shift of the energy levels $\bm{E}\rightarrow\bm{E}+\delta\bm{e}$, we have the displacement symmetry
$P_{n}(\bm{E}+\delta\bm{e})\approx P_{n}(\bm{E})+\delta\bm{e}\cdot\nabla P_{n}=P_{n}(\bm{E}),$
since $\bm{e}\cdot\nabla P_{n}=\partial_{n}(\bm{e}\cdot\nabla F)=0$.
Statistical distribution from QTI - To obtain the distribution vector $\bm{P}$ from the EIE based on QTI, we consider the symmetry of $F=F(E_{1},\cdots,E_{i},\cdots,E_{j},\cdots)$. Because the internal energy $U=\sum_{n}E_{n}P_{n}$ possesses the exchanging symmetry under operation: $(E_{i}\leftrightarrow E_{j})$, $(P_{i}\leftrightarrow P_{j})$, which guarantees the invariance of statistical distributions $(P_{1},P_{2},\cdots,P_{N})$. Up to a same scaling $\beta$ (with Eq. (8)), the EIE has a symmetric solution
$F=\left(1/\beta\right)\Psi(\beta\bm{E})=\left(1/\beta\right)\Psi(\beta E_{1},\beta E_{2},\cdots,\beta E_{N}),$ (9)
where $\Psi$ is a symmetric function of $N$ variables.
It should be noticed that the $E_{n}$ only appears in $F$ in the form of its product with $\beta$. The QTI reflects the fact that the thermodynamic entropy $S$ is a state function. We can verify that $\mathrm{d}S=\beta\delta Q$, as a total differential, satisfies $S=\beta(U-F)$, which refers to the Second Law of Thermodynamics. Following the Carathéodory’s arguments, it can be proved within our framework that the entropy of an arbitrary adiabatic process never decreases (Carathéodory, 1909; Born, 1921; Pogliani and Berberan-Santos, 2000).
We notice that the statistical distribution $P_{n}(\bm{E})=\partial_{n}\left(\Psi(\beta\bm{E})/\beta\right)$ usually depends on all energy levels. In the absence of any correlation between two “events” occupying $E_{n}$ and $E_{m}$, the relative probability $P_{n}(\bm{E})/P_{m}(\bm{E})$ does not depend on any other level $E_{l}$ ($l\neq m,n$) besides $E_{n}$ and $E_{m}$. This is consistent with the principle of detailed balance, which states that the rate of each transition process between levels $m$ and $n$ is exactly balanced by the rate of its reverse process (Breuer, 2007). In this sense, we can prove (see Sec. II of the Supplementary Materials (SM_, )) that the symmetry function $\Psi(x_{1},x_{2},\cdots,x_{N})$ reduces to $\Psi=\Psi(\sum_{n=1}^{N}w(x_{n}))$, which corresponds to
$P_{n}=\frac{w^{\prime}(E_{n})}{\sum_{m}w^{\prime}(E_{m})}.$
The EIE further determines the concrete form of $w$ as $w(E_{n})=\exp(-\beta E_{n})$, giving the usual Boltzmann distribution
$P_{n}=\frac{1}{Z}\mathrm{e}^{-\beta E_{n}},$
where $Z=\sum_{n}\exp(-\beta E_{n})$ is the partition function with respect to the free energy function $F=-k_{B}T\ln Z$.
Non-canonical states from EIE - Non-canonical statistics usually represent the equilibrium states without a well-defined temperature. Here, the detailed balance is not required for equilibrium. Instead, the inverse temperature distributes around the given integrating factor $\beta$ with a standard deviation $\Delta\beta$ (Beck and Cohen, 2003; Falcioni et al., 2011; Fei and Ma, 2024). Non-canonical statistics are usually related to systems in contact with reservoirs of finite size out of the thermodynamic limit, which is characterized by a finite heat capacity (Xu et al., 2014; Richens et al., 2018; Chung and Hassanabadi, 2019).
Interestingly, a linear superposition of free energy $F^{(\mathrm{nc})}(\beta,\bm{E})=\int\mathrm{d}\beta^{\prime}f(\beta,\beta^{\prime})F(\beta^{\prime},\boldsymbol{E})$ over varying temperatures results in a non-canonical state
$P_{n}^{(\mathrm{nc})}(\beta)=\int\mathrm{d}\beta^{\prime}f(\beta,\beta^{\prime})P_{n}(\beta^{\prime}).$ (10)
Here, the coefficient $f(\beta,\beta^{\prime})$ satisfies $\partial_{\beta}[\beta f(\beta,\beta^{\prime})]+\beta^{\prime}\partial_{\beta^{\prime}}f(\beta,\beta^{\prime})=0$ (see Sec. III of the Supplementary Materials (SM_, ) for details). Without loss of generality, we can rescale $\beta$ so that it represents the average inverse temperature. The above non-canonical distributions will show the information correlation between two events associated with $E_{m}$ and $E_{n}$.
When the standard deviation $\Delta\beta$ of $\beta$ is sufficiently small, the non-canonical states (10) are expanded as
$P_{n}^{(\mathrm{nc})}(\beta,\Delta\beta)\approx\frac{1}{Z_{\mathrm{eff}}}\exp[-\beta_{\mathrm{eff}}E_{n}+(\Delta\beta^{2}/2)E_{n}^{2}],$ (11)
where $\beta_{\mathrm{eff}}\equiv\beta+U\Delta\beta^{2}$ and $Z_{\mathrm{eff}}\equiv Z\mathrm{e}^{-\Delta\beta^{2}(U^{2}-\left\langle\Delta E^{2}\right\rangle)/2}$, and $\left\langle\Delta E^{k}\right\rangle$ is the $k$-th central moment of energy for the canonical states. With a small deviation, the detailed shape of function $f(\beta,\beta^{\prime})$ becomes inessential, where all the non-canonical thermodynamic quantities are expressed as functions of the standard deviation $\Delta\beta$ and equilibrium quantities (see Sec. IV of the Supplementary Materials (SM_, )). Therefore, without loss of generality, the coefficient $f(\beta^{\prime},\beta)$ can be chosen as a Gaussian distribution $f(\beta,\beta^{\prime})\sim\mathcal{N}(\beta^{2},\Delta\beta^{2})$ in the following calculations. It follows from the EIE that the relative deviation $\sigma\equiv\Delta\beta/\beta$ for such states remains constant.
Figure 2: Difference in the canonical and non-canonical states for three levels system where $\Delta_{1}=1,\Delta_{2}=2$. The markers are obtained with numerical calculation, and the curves depict the corresponding leading terms. (a) Comparing the distribution probabilities. The light blue bars show the canonical distribution, and the dark blue bars show the non-canonical distribution with $\beta=1,\sigma=0.5$. (b) Relative entropy between the canonical and non-canonical distributions. (c) Difference in internal energy $\Delta U$ (red circles and curve) and generalized free energy $\Delta F$ (blue squares and curve). (d) Difference in information entropy (blue squares and curve) and thermodynamic entropy (red circles and curve).
To further explore the property of the above non-canonical states, we investigate an example of three level systems with $E_{0}=0$, $E_{1}=\Delta_{1}$ and $E_{2}=\Delta_{2}$. In FIG. 2(a-b), we compare the difference in distributions in the canonical and non-canonical states. FIG. 2(a) shows the probabilities at each level for the canonical (light blue bars) and the non-canonical (dark blue bars) states where $\sigma=0.3$. FIG. 2(b) shows the relative entropy
$D_{KL}(P_{n}||P_{n}^{(\mathrm{nc})})\equiv-\sum_{n}P_{n}\ln\dfrac{P_{n}^{(\mathrm{nc})}}{P_{n}},$
between two distributions. In FIG. 2(c-d), the quantities of non-canonical states are calculated to compare with those of canonical states, which are evaluated both numerically and analytically as functions of the relative deviation $\sigma$. Respectively, FIG 2(c) illustrates the difference in the internal energy $\Delta U\equiv U^{(\mathrm{nc})}-U$ (red circles and curve) and the generalized free energy $\Delta F\equiv F^{(\mathrm{nc})}-F$ (blue squares and curve), and FIG 2(d) shows the difference in information entropy (blue squares and curve) and thermodynamic entropy (red circles and curve) for the non-canonical distributions. In the figure, all markers are evaluated numerically, and the curves represent the leading order. Here we have chosen , $\Delta_{1}=1$, $\Delta_{2}=2$, and $\beta=1$.
Correlation from Non-canonical state - For the above non-canonical states defined by Eq. (10), we can check that $S_{\mathrm{in}}(P_{n},P_{m}^{\prime})\neq S_{\mathrm{in}}(P_{m})+S_{\mathrm{in}}(P_{n})$ for $S_{\mathrm{in}}(P_{n})=-\sum_{n}P_{n}\ln P_{n}$. The information entropy for the subsystems just shows no additivity due to the non-canonicality. Such correlation is quantified by the mutual information $I\equiv S_{\mathrm{in}}(P_{m})+S_{\mathrm{in}}(P_{n})-S_{\mathrm{in}}(P_{n},P_{m}^{\prime})$. We plot the mutual information of two correlated three-level systems as a function of relative deviation in FIG. 3.
Figure 3: Mutual information between two three levels systems as a function of the relative deviation of inverse temperature.
Actually, the non-canonicality of a black hole induces the information embedded in the correlation of its radiation, which has been considered a promising candidate for the black hole information paradox. The question arises whether the non-canonical solutions to EIE can describe such non-canonical radiation field from a small Schwarzchild back hole. For a Schwarzchild black hole with temperature $\beta=8\pi M$, the deviation of $\beta$ is estimated as $\Delta\beta^{2}=\left|\beta^{2}/C\right|=8\pi$, where $C=-8\pi M^{2}$ is the heat capacity of the black hole (Hawking, 1975; Wald, 2001; Landau and Lifshitz, 2013; Carlip, 2014). From non-canonical solutions of EIE, the emission rate (see Sec. V of Supplementary Materials (SM_, )) reads as
$\Gamma(\beta,\omega)\approx\frac{1}{4\pi^{2}}\exp[-8\pi\omega(M-\omega/2)],$ (12)
which is valid in the cases with small deviation. This result is in agreement with the former results obtained from the microscopic theory for the radiation field near the small black hole (Parikh and Wilczek, 2000). The non-canonical state is capable of accounting for the black hole information paradox (Zhang et al., 2009; Cai et al., 2016). It is emphasized that such a non-canonical solution in this letter is regarded as the result of thermodynamic integrability. As black holes become smaller, the back-reaction of radiation grows increasingly stronger, leading to a tighter correlation between successive radiation events. This heightened correlation causes the radiation state to deviate significantly from the traditional canonical thermal equilibrium state.
Conclusion - In conclusion, we have proposed an alternative approach to reconsider the foundations of statistical mechanics according to the quantum thermodynamic integrability (QTI) that we introduced in this letter as a generalization of the path-independence of work and heat. We established a quantum version of Carathéodory’s principle and derived the so-called entropy integrable equations (EIE), which can result in both canonical and non-canonical states. By considering detailed balance, we uniquely constructed the Boltzmann distribution from EIE for QTI as the foundation of statistical physics. In this sense, QTI not only reinforces the principles of statistical physics based on the experiment-driven principles of thermodynamics, but also provides a new perspective on non-canonical thermodynamics based on information correlation. This work opens up new avenues for addressing longstanding problems in statistical thermodynamics, offering a more comprehensive framework for analyzing complex thermodynamic behaviors in small systems.
Acknowledgment - We are grateful to H. Dong, H. T. Quan and Y. H. Ma for their helpful discussions. This work has been supported by the National Natural Science Foundation of China (NSFC) (Grant No. 12088101) and NSAF No.U2330401.